# Breaking Focus: Contextual Distraction Curse in Large Language Models

### v1

@@ Abstract
Abstract: Recent advances in Large Language Models (LLMs) have revolutionized generative systems, achieving excellent performance across diverse domains. Although these models perform well in controlled environments, their real-world applications frequently encounter inputs containing both essential and irrelevant details. Our investigation has revealed a critical vulnerability in LLMs, which we term Contextual Distraction Vulnerability (CDV). This phenomenon arises when models fail to maintain consistent performance on questions modified with semantically coherent but irrelevant context. To systematically investigate this vulnerability, we propose an efficient tree-based search methodology to automatically generate CDV examples. Our approach successfully generates CDV examples across four datasets, causing an average performance degradation of approximately 45% in state-of-the-art LLMs. To address this critical issue, we explore various mitigation strategies and find that post-targeted training approaches can effectively enhance model robustness against contextual distractions. Our findings highlight the fundamental nature of CDV as an ability-level challenge rather than a knowledge-level issue since models demonstrate the necessary knowledge by answering correctly in the absence of distractions. This calls the community’s attention to address CDV during model development to ensure reliability. Code is available at https://github.com/wyf23187/LLM_CDV.
@@ 1 Introduction
Large Language Models (LLMs) (Zhou et al., 2023b) have demonstrated remarkable capabilities across various Natural Language Processing (NLP) tasks, revolutionizing wide downstream applications such as medicine (Zhao et al., 2023), education (Kasneci et al., 2023), and science (Li et al., 2024b; Guo et al., 2023; Huang et al., 2024e). Despite their impressive performance, recent studies have exposed various vulnerabilities in LLMs, including susceptibility to jailbreaking attacks (Zou et al., 2023), hallucination issues (Xu et al., 2024b), and consistency problems (Liang et al., 2024; Huang et al., 2024a). These vulnerabilities highlight the limitations of LLMs in handling nuanced and adversarial scenarios, making it critical to uncover and analyze additional weaknesses to improve their reliability.
In this work, we investigate a novel vulnerability termed Contextual Distraction Vulnerability (CDV), where semantically coherent but non-essential contextual additions to a question degrade LLM performance. For instance, a customer service chatbot might miss a refund request hidden in a short story about discovering products through social media influencers. Similarly, a technical query about machine learning could be misunderstood if it’s preceded by a student’s emotional account of exam preparation anxiety. Unlike adversarial attacks that inject semantically meaningless noise into inputs (Zou et al., 2023; Shi et al., 2024) and distraction brought by long-context input (Bai et al., 2023), for CDV, our study demonstrates that semantically coherent without a long context yet contextually distracting modifications are sufficient to disrupt the decision-making process of even the most advanced LLMs. This vulnerability underscores a critical weakness in LLMs’ ability to filter out irrelevant information and prioritize core knowledge, which is essential for robust reasoning.
Recent studies have demonstrated the powerful generative capabilities of LLM Xu et al. (2024a); Wu et al. (2024), To systematically investigate this vulnerability, we propose a methodology for efficiently automating the generation of questions that may trigger CDV in LLMs. This is a non-trivial problem, as the primary challenge lies in identifying the most effective distraction (a perturbation to the original input) within a vast search space while ensuring that the added distraction remains semantically coherent with the original question. Our proposed method employs a tree-based search algorithm, incorporates early stopping strategies and leverages a simple but effective classifier to pre-filter problematic candidates. Through extensive experimentation across multiple datasets on the latest LLMs, we demonstrate the effectiveness of our method in producing effective CDV question examples.
Beyond this, we investigate potential causes and mitigation strategies for CDV. Our findings reveal that 1) while simple prompt engineering provides limited benefits, targeted fine-tuning (e.g., DPO (Rafailov et al., 2024)) significantly enhances model robustness against contextual distractions. 2) CDV is a fundamental challenge in LLM development that requires ability-level rather than knowledge-level enhancement.
In summary, our contributions are three-folds: 1) We identify and characterize CDV, a significant weakness in current LLMs that affects their reliability in question-answering tasks. 2) We propose an efficient automated pipeline to generate semantically valid contextual distractions that expose this vulnerability, achieving substantial performance degradation across various models and datasets. 3) We conduct comprehensive experiments to evaluate different mitigation strategies, providing insights into the nature of CDV and potential approaches to overcome CDV.
2 Related Work
LLM Vulnerability. Previous research has extensively explored various vulnerabilities in LLMs. Jailbreak attacks (Wei et al., 2024; Zou et al., 2023; Huang et al., 2024d) have highlighted security risks, driving efforts to develop more trustworthy LLMs (Huang et al., 2024c; Liu et al., 2023; Huang et al., 2023b). Hallucinations (Li et al., 2023) remains a persistent issue in various scenarios like LLM-based agents (Zhang et al., 2024b; Huang et al., 2023a), leading to incorrect or misleading responses Huang et al. (2024b). Additionally, LLMs are highly susceptible to prompt injection attacks (Liu et al., 2024), where adversarial prompts disguise malicious instructions as benign inputs. Berglund et al. (2023) identify the reversal curse in LLMs, exposing their failure to generalize in auto-regressive settings. Sharma et al. (2023) investigate sycophancy, revealing that human feedback may inadvertently encourage LLMs to align with user beliefs rather than provide truthful responses. Furthermore, studies on LLM honesty (Yang et al., 2023; Gao et al., 2024) suggest that LLMs often lack self-awareness regarding their capabilities, contributing to hallucinations and unreliable outputs.
Our study focuses on a novel vulnerability–CDV, in which LLMs lose focus due to contextual distractions. We identify CDV as a fundamental ability-level challenge for LLMs, parallel to the aspects previously explored. Addressing CDV is therefore crucial for enhancing LLMs’ reliability alongside the already assessed capabilities.
Figure 1: The overview of the proposed method. Given a problem, our goal is to automatically transform it into a CDV example. Initially, a classifier filters the problems to identify potential candidates that are easy to perturb. Next, the method employs a tree-based search, supported by a priority queue to manage the search priorities of individual nodes. Using Error-Guided Perturbation (EGP) generation, the pipeline efficiently and automatically produces effective CDV examples.
Tree-Based Search for LLMs. Tree search methods have shown considerable potential in enhancing the exploratory capabilities of LLMs (Zhang et al., 2023; Hu et al., 2024). Tree Prompting (Singh et al., 2023) offers a novel approach that solves problems by constructing decision prompt trees. Zhou et al. (2023a) and Koh et al. (2024) treat the language model as a general agent, utilizing tree-based search to enhance the model capability of both reasoning and planning. Building upon the Chain-of-Thought (CoT) (Wei et al., 2022), Tree of Thoughts (ToT) (Yao et al., 2024) extends this idea by decomposing reasoning tasks into sequential steps and exploring potential thoughts at each stage. Moreover, Li et al. (2024a) proposes CodeTree to efficiently explore the search space in different stages during code generation.
Our proposed method employs a tree-based search to systematically explore potential perturbations of the original input, aiming to automatically generate CDV samples that effectively challenge LLMs.
@@ 3.1 Overview
As illustrated in Figure 1, our objective is to systematically identify and analyze the vulnerabilities of LLMs by generating high-quality CDV examples that challenge models while maintaining the semantic integrity of the original questions.
Generating effective CDV examples poses three significant challenges: First, it is critical to design an effective reward signal that guides modifications toward high-quality CDV examples. Second, the process must address the computational overhead caused by the vast search space associated with generating and evaluating these examples. Third, the generation must mitigate semantic shifts and account for the impact introduced by long-context inputs to ensure the original intent of the examples.
To overcome these challenges, we introduce a multi-step framework. Initially, a classifier-based filtering mechanism is employed to identify original samples that are easy to perturb, substantially reducing the computational burden by narrowing down the search space. For the filtered samples, we apply a tree-based search strategy that systematically explores contextual distractions using simulation-driven evaluation. This process generates controlled perturbations through a proxy model and guides the search using rewards based on the outcomes of the victim model during simulation. Iterative semantic validation and length control are applied to ensure that the generated examples align closely with the original intent. Furthermore, we incorporate efficient pruning and early stopping techniques to balance the trade-off between uncovering meaningful CDV examples and maintaining computational efficiency.
@@ 3.2 Problem Formulation
with $Q$ denoting the question, $A_{\text{gt}}$ the ground truth answer, and $\mathcal{A}_{\text{inc}}$ a set of incorrect answers. Given a victim model $M$, our goal is to construct a perturbed dataset $D^{\prime}=\{P^{\prime}_{1},P^{\prime}_{2},\dots,P^{\prime}_{N}\}$, where each perturbed instance $P^{\prime}=\langle Q^{\prime},A_{\text{gt}},\mathcal{A}_{\text{inc}}\rangle$ is obtained by applying a perturbation $\Delta Q$ to the question $Q$, such that:
@@
Our aim is to optimize the perturbation $\Delta Q$ to minimize the accuracy of $M$ on $D^{\prime}$, while ensuring semantic consistency and length constraints between $Q$ and $Q^{\prime}$. Here, semantic consistency is determined by a binary classifier $S$, which outputs $S(Q,Q^{\prime})\in\{0,1\}$, where $S(Q,Q^{\prime})=1$ indicates no semantic shift. Formally, the problem is expressed as:
$\min_{\Delta Q}$ $\quad\mathbb{E}_{P\sim D}\Big{[}\mathcal{L}_{\text{accuracy}}(M,Q^{\prime})\Big{]},$ (3)
s.t. $\quad S(Q,Q^{\prime})=1,\quad\frac{\text{len}(Q^{\prime})}{\text{len}(Q)}\leq\lambda,$
Here, $S$ ensures that the perturbation $\Delta Q$ does not lead to a semantic shift, while the length ratio constraint $\lambda$ ensures $Q^{\prime}$ remains within acceptable bounds compared to the original $Q$. This constraint is necessary because recent studies show that LLMs experience performance degradation in long context scenarios (Bai et al., 2023). To prevent excessive length expansion in $Q^{\prime}$, we introduce a length constraint, where $\lambda$ is an upper bound on the relative length of $Q^{\prime}$ compared to $Q$.
If the output $Q^{\prime}$ does not satisfy the constraints in Equation 3, it is discarded, and a new perturbation $\Delta Q$ is generated by re-prompting the proxy model.
@@ 3.3 Error-Guided Perturbation Generation
The perturbation $\Delta Q$ is generated using a proxy model, denoted as $P_{\text{proxy}}$. The proxy model is prompted with the original problem instance $P=\langle Q,A_{\text{gt}},\mathcal{A}_{\text{inc}}\rangle$ and tasked with generating a modified question $Q^{\prime}$ defined in Equation 2, where $\Delta Q$ represents the perturbation introduced by $P_{\text{proxy}}$. The perturbation process is formalized as:
@@
where $P_{\text{proxy}}$ generates $\Delta Q$ based on a predefined prompt designed to guide the proxy model in producing modifications that increase the likelihood of the victim model $M$ selecting an incorrect answer $a_{\text{inc}}\in\mathcal{A}_{\text{inc}}$ (i.e., lead model to response with an error).
@@ 3.4 Tree-Based Perturbation Exploration
We employ a tree-based simulation-driven method to optimize perturbations by heuristically exploring the search space. A priority queue is maintained to store nodes ordered by their value $\mathcal{V}(P^{\prime})$, with the highest-value node dequeued and expanded iteratively using $P_{\text{proxy}}$ to identify high-potential vulnerabilities.
Simulation For Measuring Perturbation Quality. Firstly, we aim to design the reward of perturbed questions to measure their value. For a given problem instance $P=\langle Q,A_{\text{gt}},\mathcal{A}_{\text{inc}}\rangle$, the simulation process evaluates the quality of a perturbation by estimating the success rate of the victim model $M$ on $P$. Let $y\sim M(y\mid P)$ represent the output of the model $M$ when queried on $P$. During a single simulation, the success rate $r_{M}(P)$ is computed by sampling $n$ model outputs:
@@
where $\mathbb{I}\{\cdot\}$ is an indicator function that returns 1 if the model’s output $y_{i}$ matches the ground truth answer $A_{\text{gt}}$, and 0 otherwise. The success rate $r_{M}(P)$ quantifies the likelihood of the model producing the correct answer under the given perturbation.
A perturbed problem $P^{\prime}=\langle Q^{\prime},A_{\text{gt}},\mathcal{A}_{\text{inc}}\rangle$ is considered successfully perturbed if $r_{M}(P^{\prime})=0$, indicating that the model fails to produce the correct answer in all sampled outputs. The simulation process computes a value $\mathcal{V}(P^{\prime})$ for the node corresponding to $P^{\prime}$ in the tree-based search:
@@
where $\alpha$ and $\gamma$ are scaling constants, $r_{M}(P^{\prime})$ is the success rate of the victim model $M$ on $P^{\prime}$, and depth represents the recursion depth of the node in the search tree. For the question with $r_{M}(P^{\prime})=0$, we add it into the candidate problem list $L$ for storing the problems considered to be successfully “frustrate” LLMs.
The simulation process systematically estimates $\mathcal{V}(P^{\prime})$, prioritizing perturbed problems with lower success rates $r_{M}(P^{\prime})$, which correspond to higher potential vulnerabilities in the model. Simultaneously, the factor $\text{depth}^{-\gamma}$ discourages deeper recursions in the search tree, ensuring computational efficiency. High-value nodes with large $\mathcal{V}(P^{\prime})$ scores are prioritized in the following tree-based search.
@@
The proxy model $P_{\text{proxy}}$ generates $k=|\mathcal{A}_{\text{inc}}|$ child nodes for $P^{\prime}$, corresponding to perturbations $\Delta Q_{j}$ derived from each incorrect candidate answer $a_{\text{inc}}\in\mathcal{A}_{\text{inc}}$:
@@
The search iteratively repeats this searching process, dynamically expanding the highest-value node and exploring the perturbation space.
Why not Monte Carlo Tree Search? Monte Carlo Tree Search (MCTS) (Browne et al., 2012) has been widely used in recent studies to perform simulations powered by LLMs, achieving remarkable performance (Zhang et al., 2024a; Wang et al., 2024; Xie et al., 2024; Guan et al., 2025). However, MCTS is not suitable for our task due to its focus on balancing exploration (searching broadly across the tree) and exploitation (focusing on promising branches). In our context, such a balance is unnecessary because the width of the tree is inherently fixed, dictated by the number of incorrect answer candidates $|\mathcal{A}_{\text{inc}}|$. Moreover, MCTS introduces computational overhead by maintaining dynamic exploration strategies, which is impractical given the predefined structure and requirements of our method. Therefore, we opt for a simpler and more task-specific tree design that aligns directly with the properties of our problem.

### v2

# Adaptive Distraction: Probing LLM Contextual Robustness with Automated Tree Search

@@ Abstract
Large Language Models (LLMs) often struggle to maintain their original performance when faced with semantically coherent but task-irrelevant contextual information. Although prior studies have explored this issue using fixed-template or retrieval-based distractions, such static methods show limited effectiveness against contemporary models. To address this problem, we propose a dynamic distraction generation framework based on tree search, where the generation process is guided by model behavior. Without modifying the original question or answer, the method efficiently produces challenging adaptive distractions across multiple datasets, enabling systematic stress testing of LLMs’ contextual robustness. Experiments on four benchmarks demonstrate that the generated distractions lead to an average performance drop of over 45% for mainstream models. Further comparisons of mitigation strategies show that prompt-based optimization methods yield limited gains, whereas post-training approaches (e.g., DPO) significantly enhance the model’s contextual robustness. The results indicate that these issues do not stem from knowledge deficits in LLMs, but from a fundamental inability to maintain consistent reasoning under contextual distraction, posing a major challenge to the reliability of LLMs in real-world applications. The code is publicly available at https://github.com/wyf23187/Adaptive_Distractions.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success across diverse natural language processing tasks, such as question answering, summarization, and reasoning achiam2023gpt ; zhou2023survey ; huang2025trustgen ; xu2025socialmaze . However, recent studies reveal a critical vulnerability: LLMs are susceptible to semantically coherent but task-irrelevant contextual information, which can significantly degrade their performance shi2023large . This lack of contextual robustness hinders their ability to consistently focus on essential task content in the presence of distracting information, a challenge particularly pronounced in real-world applications where irrelevant context is common. Addressing this limitation is crucial to ensure the reliability of LLMs in complex, dynamic environments.
Current methods for evaluating LLM contextual robustness primarily rely on fixed-template or retrieval-based distractors shi2023large ; liu2023lost . However, our preliminary experiments demonstrate that these static approaches are increasingly ineffective against contemporary models, with performance degradation often below 5% on advanced models like GPT-4o, which is detailed in Appendix B.2. This highlights that existing methods lack adaptiveness and are heavily dependent on the specific behavior of the target LLM. Once the model evolves, previously effective attack strategies may become obsolete. Moreover, such limited impact is insufficient to provide a reliable basis for robustness evaluation in realistic scenarios shayegani2023survey . This highlights the urgent need for dynamic, adaptive methods capable of generating contextually plausible distractions that evolve with LLM capabilities, ensuring robust evaluation across diverse tasks and models.
To address the limitations of static methods, we aim to propose an adaptive attack method to generate adaptive distractions—semantically coherent, answer-preserving contextual additions that significantly impair LLM performance. This approach aims to evolve with advancing LLM capabilities, enabling robust stress-testing across diverse tasks without being constrained by model strength. However, generating such distractions presents key challenges: (1) ensuring semantic coherence with the original input, (2) preserving the correct answer, and (3) creating distractions potent enough to disrupt model predictions. These requirements demand a dynamic, model-informed generation strategy to effectively probe LLM contextual robustness.
To address these challenges, we propose a structured generation framework based on tree search to automatically construct adaptive distractions yao2023tree ; browne2012survey ; xu2025cross . Our approach employs a classifier to pre-filter questions susceptible to perturbation, followed by a tree search module that explores contextual additions using a priority queue guided by model behavior. Error-guided perturbations generate candidate distractions at each node, evaluated for their ability to alter predictions while preserving semantic consistency. Early stopping strategies enhance efficiency, ensuring scalability across tasks and model families. Figure 1 illustrates this pipeline, which enables the controlled and automated creation of challenging distractions tailored to probe LLM contextual robustness.
We conducted comprehensive experiments to validate our framework, evaluating its effectiveness across four benchmark datasets, namely MMLU, CommonsenseQA, OpenbookQA, and TruthfulQA, on a diverse set of mainstream models, including proprietary and open-weight architectures. Our results show that adaptive distractions cause significant performance degradation, with an average accuracy drop exceeding 45%, exposing vulnerabilities in even the most advanced LLMs. Additionally, we explored mitigation strategies, comparing prompt-based approaches with targeted fine-tuning, and analyzed supplementary experiments, including prompt variants and case studies, detailed in the appendix. These findings collectively underscore the critical need for enhanced contextual robustness in LLMs.
In summary, our work delivers the following contributions:
  • We introduce a novel framework for generating adaptive distractions, semantically coherent yet task-irrelevant additions, enabling robust and systematic evaluation of LLM contextual robustness.
  • We provide empirical evidence of significant performance degradation, exceeding 45% accuracy drop, across four benchmark datasets and diverse mainstream models, uncovering persistent vulnerabilities in advanced LLMs.
  • We evaluate mitigation strategies, revealing that targeted fine-tuning substantially enhances robustness under contextual distraction, while prompt-based approaches yield limited effectiveness.
@@ 2.1 Overview
As illustrated in Figure 1, our objective is to systematically identify vulnerabilities in LLMs by generating adaptive distractions. As defined in the Introduction, these are contextual additions designed to preserve the original question’s meaning and answer while affecting model performance.
Figure 1: Overview of our framework. Given an input question, a classifier first filters for instances that are more susceptible to contextual interference. Then, a tree-based search explores candidate context additions using error-guided perturbation (EGP). At each node, candidate distractions are evaluated based on their ability to alter model predictions without affecting the correct answer. The framework efficiently produces high-quality adaptive distraction examples that challenge model robustness.
Generating such distractions presents three core challenges. First, it requires an effective reward mechanism to guide the generation toward minimal yet impactful context additions. Second, the search space for valid distractions is vast, leading to substantial computational cost if not carefully controlled. Third, ensuring semantic consistency and bounded input length is critical, as irrelevant context may introduce unintended shifts or exceed the model’s optimal context window.
To address these challenges, we propose a multi-step framework. We first apply a classifier-based filtering process to identify examples that are more likely to be affected by contextual interference, which narrows the search space. Then, we perform a tree-based search to explore semantically valid distractions. At each node, a proxy model generates candidate perturbations, which are evaluated through simulation using the victim model. The search is guided by a value function that combines model failure signals with depth penalties to ensure both quality and efficiency. Finally, we incorporate semantic validation and length control to preserve the original problem’s intent, alongside pruning and early stopping strategies to reduce computation without sacrificing attack effectiveness.
@@ 2.2 Problem Formulation
with $Q$ denoting the question, $A_{\text{gt}}$ the ground truth answer, and $\mathcal{A}_{\text{inc}}$ a set of incorrect answers. Given a victim model $M$, our goal is to construct a perturbed dataset $D^{\prime}=\{P^{\prime}_{1},P^{\prime}_{2},\dots,P^{\prime}_{N}\}$, where each perturbed instance $P^{\prime}=\langle Q^{\prime},A_{\text{gt}},\mathcal{A}_{\text{inc}}\rangle$ is obtained by applying an adaptive distraction $\Delta Q$ to the question $Q$, such that:
@@
Our aim is to optimize the distraction $\Delta Q$ to minimize the accuracy of $M$ on $D^{\prime}$, while ensuring semantic consistency and length constraints between $Q$ and $Q^{\prime}$. Here, semantic consistency is determined by a binary classifier $S$, which outputs $S(Q,Q^{\prime})\in\{0,1\}$, where $S(Q,Q^{\prime})=1$ indicates no semantic shift. Formally, the problem is expressed as:
$\min_{\Delta Q}\quad\mathbb{E}_{P\sim D}\Big[\mathcal{L}_{\text{accuracy}}(M,Q^{\prime})\Big],\text{s.t.}\quad S(Q,Q^{\prime})=1,\quad\frac{\text{len}(Q^{\prime})}{\text{len}(Q)}\leq\lambda,$ (3)
Here, $S$ ensures that the distraction $\Delta Q$ does not lead to a semantic shift, while the length ratio constraint $\lambda$ ensures $Q^{\prime}$ remains within acceptable bounds compared to the original $Q$. This constraint is necessary because recent studies show that LLMs experience performance degradation in long context scenarios bai2023longbench . To prevent excessive length expansion in $Q^{\prime}$, we introduce a length constraint, where $\lambda$ is an upper bound on the relative length of $Q^{\prime}$ compared to $Q$.
If the output $Q^{\prime}$ does not satisfy the constraints in Equation 3, it is discarded, and a new distraction $\Delta Q$ is generated by re-prompting the proxy model.
@@ 2.3 Error-Guided Perturbation Generation
The distraction $\Delta Q$ is generated using a proxy model, denoted as $P_{\text{proxy}}$. The proxy model is prompted with the original problem instance $P=\langle Q,A_{\text{gt}},\mathcal{A}_{\text{inc}}\rangle$ and tasked with generating a modified question $Q^{\prime}$ defined in Equation 2, where $\Delta Q$ represents the distraction introduced by $P_{\text{proxy}}$. The generation process is formalized as:
@@
where $P_{\text{proxy}}$ generates $\Delta Q$ based on a predefined prompt designed to guide the proxy model in producing contextual additions that increase the likelihood of the victim model $M$ selecting an incorrect answer $a_{\text{inc}}\in\mathcal{A}_{\text{inc}}$ (i.e., lead the model to produce an error).
@@ 2.4 Tree-Based Perturbation Exploration
We employ a tree-based simulation-driven method to optimize distractions by heuristically exploring the search space. A priority queue is maintained to store nodes ordered by their value $\mathcal{V}(P^{\prime})$, with the highest-value node dequeued and expanded iteratively using $P_{\text{proxy}}$ to identify high-potential vulnerabilities.
Simulation For Measuring Distraction Quality. Firstly, we aim to design the reward of distracted questions to measure their value. For a given problem instance $P=\langle Q,A_{\text{gt}},\mathcal{A}_{\text{inc}}\rangle$, the simulation process evaluates the quality of a distraction by estimating the success rate of the victim model $M$ on $P$. Let $y\sim M(y\mid P)$ represent the output of the model $M$ when queried on $P$. During a single simulation, the success rate $r_{M}(P)$ is computed by sampling $n$ model outputs:
@@
where $\mathbb{I}\{\cdot\}$ is an indicator function that returns 1 if the model’s output $y_{i}$ matches the ground truth answer $A_{\text{gt}}$, and 0 otherwise. The success rate $r_{M}(P)$ quantifies the likelihood of the model producing the correct answer under the given distraction.
A distracted problem $P^{\prime}=\langle Q^{\prime},A_{\text{gt}},\mathcal{A}_{\text{inc}}\rangle$ is considered effective if $r_{M}(P^{\prime})=0$, indicating that the model fails to produce the correct answer in all sampled outputs. The simulation process computes a value $\mathcal{V}(P^{\prime})$ for the node corresponding to $P^{\prime}$ in the tree-based search:
@@
where $\alpha$ and $\gamma$ are scaling constants, $r_{M}(P^{\prime})$ is the success rate of the victim model $M$ on $P^{\prime}$, and depth represents the recursion depth of the node in the search tree. For the question with $r_{M}(P^{\prime})=0$, we add it into the candidate problem list $L$, which stores examples that effectively induce failure in the LLM.
The simulation process systematically estimates $\mathcal{V}(P^{\prime})$, prioritizing distracted problems with lower success rates $r_{M}(P^{\prime})$, which correspond to higher potential vulnerabilities in the model. Simultaneously, the factor $\text{depth}^{-\gamma}$ discourages deeper recursions in the search tree, ensuring computational efficiency. High-value nodes with large $\mathcal{V}(P^{\prime})$ scores are prioritized in the following tree-based search.
@@
The proxy model $P_{\text{proxy}}$ generates $k=|\mathcal{A}_{\text{inc}}|$ child nodes for $P^{\prime}$, corresponding to distractions $\Delta Q_{j}$ derived from each incorrect candidate answer $a_{\text{inc}}\in\mathcal{A}_{\text{inc}}$:
@@
The search iteratively repeats this searching process, dynamically expanding the highest-value node and exploring the distraction space.
Why not Monte Carlo Tree Search? Monte Carlo Tree Search (MCTS) browne2012survey has been widely used in recent studies to perform simulations powered by LLMs, achieving remarkable performance zhang2024rest ; wang2024seed ; xie2024monte ; guan2025rstar . However, MCTS is not suitable for our task due to its focus on balancing exploration (searching broadly across the tree) and exploitation (focusing on promising branches). In our context, such a balance is unnecessary because the width of the tree is inherently fixed, dictated by the number of incorrect answer candidates $|\mathcal{A}_{\text{inc}}|$. Moreover, MCTS introduces computational overhead by maintaining dynamic exploration strategies, which is impractical given the predefined structure and requirements of our method. Therefore, we opt for a simpler and more task-specific tree design that aligns directly with the properties of our problem.